OpenStudy (anonymous):

What is the circumference of a circle with an area of 81 in2?

OpenStudy (saifoo.khan):

find the radius first.

OpenStudy (anonymous):

circumference is 2pi(r) and 2pi(r)^2 is the area. so to find r just take 2pi(r)^2 = 81 and solve for r and then plug that into the equation of the circumference

OpenStudy (anonymous):

i still dont see it

OpenStudy (anonymous):

they give you the area and already and you need to convert it into a circumference. looking at the equation of the circumference of a circle:\[2\pi r\] and area \[2\pi r^2\]you can see that they are related by r

OpenStudy (anonymous):

right.. so i dont understand what i need to do now

OpenStudy (anonymous):

in the problem they already tell you that the area is 81

OpenStudy (anonymous):

so they already give you \[\pi r^2 = 81\]

OpenStudy (anonymous):

oh ok so the radius is 9?

OpenStudy (anonymous):

then to solve for r, you divide both sides by pi to get\[\pi r^2=81/ \pi\]

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

\[ r^2 = 81/ \pi\]

OpenStudy (anonymous):

you need to watch you algebra and follow the order of operations ->PEMDAS

OpenStudy (anonymous):

then take the square root of both sides to get\[r=\sqrt{81/ \pi}\]

OpenStudy (anonymous):

remember that pi is a number (like 3.1415 something) so its treated like it is\[ (3.14)\times \pi = 81\]

OpenStudy (anonymous):

you can check if your r is correct by plugging it into the equation of the area and you should get 81 then by the equation of the circumference and the r that we found we get that the circumference is \[2 \pi \sqrt{81/\pi} \] or if you simplify \[18 \pi^{1/2}\]

OpenStudy (anonymous):

i whoops i mean (3.14) x r = 81 not (3.14) x pi = 81

OpenStudy (anonymous):

Circumference is \[2\pi*r\] Area is \[\pi*r^2\] So the question gives you that the Area of the Circle is 81. Plug 81 into the Area formula. \[81=\pi*r^2\] From this equation you can solve and get r. So do the math and r=5.07 Now we have the value of the radius r=5.07 We now can solve for the circumference \[Circumference=2\pi*r\] \[Circumference=2\pi*(5.07)\] Do the math and your answer is 31.85 in^2